OpenStudy (anonymous):

Help!?!?!?!? 5. For the function g(x) = 14/x + 3 , find (g^–1∘ g)(4). (1 point) 6 10 4 0

OpenStudy (anonymous):

5. For the function g(x) = 14/x + 3, find (g–1∘ g)(4). (1 point) 6 10 4 0

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

g(x) is \[\Large g(x) = \frac{14}{x}+3\] OR \[\Large g(x) = \frac{14}{x+3}\] ??

OpenStudy (anonymous):

The second one :)

jimthompson5910 (jim_thompson5910):

first find the inverse \[\Large g(x) = \frac{14}{x+3}\] \[\Large y = \frac{14}{x+3}\] \[\Large x = \frac{14}{y+3}\] \[\Large x(y+3) = 14\] \[\Large xy+3x = 14\] \[\Large xy = 14-3x\] \[\Large y = \frac{14-3x}{x}\] \[\Large g^{-1}(x) = \frac{14-3x}{x}\]

jimthompson5910 (jim_thompson5910):

Now plug x = 4 into g(x). Tell me what you get.

OpenStudy (anonymous):

g^-1(4) = 2/4 or 0.5

jimthompson5910 (jim_thompson5910):

no g(x) first

jimthompson5910 (jim_thompson5910):

\[\Large g(x) = \frac{14}{x+3}\] \[\Large g(4) = \frac{14}{4+3}\] \[\Large g(4) = ?\]

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

then plug that result into the inverse

OpenStudy (anonymous):

Sorry, is the inverse the original equation?

jimthompson5910 (jim_thompson5910):

no, the original is the original i showed you how to find the inverse above

OpenStudy (anonymous):

Right! g^-1(2) = 4

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Ok cool!! Thanks for the help!!! :)

jimthompson5910 (jim_thompson5910):

yw